# SLR: Automated Synthesis for Scalable Logical Reasoning

### v4

@@ Abstract
Lukas Helff1,2, Ahmad Omar1, Felix Friedrich1,2, Antonia Wüst1, Hikaru Shindo1, Tim Woydt1, Rupert Mitchell1,2, Patrick Schramowski1,2,3,4, Wolfgang Stammer1,2,5, Kristian Kersting1,2,3,5,6 1TU Darmstadt 2hessian.AI 3DFKI 4CERTAIN, Germany 5Lab1141 6Centre for Cognitive Science, Darmstadt Code: https://github.com/ml-research/ScalableLogicalReasoning Data: https://huggingface.co/datasets/AIML-TUDA/SLR-Bench
@@ 1 Introduction
To tackle these challenges, this paper introduces SLR (Scalable Logical Reasoning), an open-source framework for evaluating and training models in inductive logical reasoning. Based on a user-defined logic task (Fig. 1, left), the task synthesizer (center) automatically generates novel inductive logic programming (ILP) tasks ILP_Muggleton; ILP_Cropper of controllable complexity. Each task comes with (i) a latent ground-truth rule, (ii) an executable validation program, and (iii) an instruction prompt for the reasoning task. The ground-truth rule serves as the reference answer, while the validation program deterministically evaluates any candidate hypothesis. SLR supports both systematic model evaluation (Fig. 1, top right) and downstream model training, via supervised finetuning or reinforcement learning with rewards provided by the integrated symbolic judge (Fig. 1, bottom right). SLR ’s fully symbolic and automated pipeline eliminates the need for human annotation and avoids dataset overlap.
Leveraging SLR, we present SLR-Bench (Fig. 3), a 19k task benchmark that forms a twenty-level curriculum of increasing logical complexity. These levels are further organized into four curriculum tiers: basic, easy, medium, and hard. Each task is unique, with a precise and systematic assessment of inductive logical reasoning skills. In evaluations, we find that while LLMs are generally well-versed in generating syntactically valid rules, robust logical reasoning remains challenging. Performance declines sharply as task complexity increases. Scaling model size brings only marginal improvements, while scaling test-time compute boosts reasoning, but returns diminish as complexity rises.
Beyond evaluation, SLR enables curriculum learning, boosting reasoning both in-domain and across established reasoning benchmarks. SLR-tuned Llama-3-8B not only surpasses all conventional LLMs on SLR-Bench, but also outperforms recent reasoning LLMs such as Gemini-2.0-flash-thinking, while using fewer inference tokens. Notably, these enhanced reasoning capabilities generalize downstream, improving performance on MMLU, LogicQA, HellaSwag, and GPQA.
@@
⟦t!⟧ Reasoning Data Evaluation Task Custom Curriculum Scalable Dataset Type Creation Methodology Synthesis Tasks Learning Complexity LogiQA liu2020logiqa Deduction Human MC LogiQA 2.0 Hanmeng2023logicqa2 Mix Human MCQA/Auto FOLIO han2022folio Deduction Human EM AbductionRules young2022abductionrules Abduction Synthetic/Human EM FineLogic zhou2025FineLogic Deduction DS collection Symbolic/LLM HLE phan2025humanitysexam Mix DS collection LLM Big-Bench phan2025humanitysexam Mix DS collection Mix LogiGLUE luo2024logiglue Mix DS collection MCQA/Auto CLUTRR sinha2019clutrr Induction Synthetic EM KOR-Bench ma2024korbenchbenchmarkinglanguagemodels Mix Synthetic/Human EM/Symbolic (✓) PrOntoQA PrOntoQA Deduction Synthetic EM SynLogic liu2025synlogic Deduction Synthetic EM FLD morishita202FLD Deduction Synthetic Symbolic (✓) K&K xie2025logicrlunleashingllmreasoning; xie2024memorization Deduction Synthetic EM (✓) (✓) ZebraLogic zebralogic2025 Deduction Synthetic EM (✓) (✓) SLR (ours) Induction Synthetic Symbolic
@@
2 Related Work
Evaluating LLMs’ Logical Reasoning. Tab. 1 provides an overview of existing logical reasoning benchmarks in terms of inference types, dataset origins, and evaluation formats. Notable datasets include LogiQA/2.0 liu2020logiqa; Hanmeng2023logicqa2, FOLIO han2022folio (deductive reasoning), AbductionRules young2022abductionrules (abductive reasoning), bAbI Weston2015TowardsAQ, and CLUTRR sinha2019clutrr (synthetic QA with inductive reasoning). Aggregate testbeds such as BIG-Bench kazemi2025bigbenchextrahard; suzgun2023bigbench, HLE phan2025humanitysexam, FineLogic zhou2025FineLogic, and LogiGLUE luo2024logiglue span a range of tasks and inference styles. Other benchmarks, such as Proofwriter, PrOntoQA, KOR-Bench, FLD, Multi-LogiEval, SynLogic, ZebraLogic, and the K&K Sandbox tafjord2021proofwriter; PrOntoQA; morishita202FLD; patel2024multilogieval; liu2025synlogic; zebralogic2025; xie2025logicrlunleashingllmreasoning; xie2024memorization; mondorf2024liar generate tasks using various logics or ontologies, often with fixed configurations or evaluation via exact match. However, most existing benchmarks lack key features such as scalability, extensibility, controlled curricula, and flexible task synthesis, all of which are addressed by SLR.
Limits and Promises of Reasoning LLMs. LLMs like GPT-4 openai2023gpt4, Llama-3 grattafiori2024llama3herdmodels, and Qwen bai2023qwen can handle basic reasoning and coding tasks but often struggle with true abstraction shojaee2025illusion; xie2024memorization. Recent reasoning LLMs attempt to bridge this gap by scaling test-time compute. Systems like OpenAI’s o1/o3 openai_o3_system_card_2025 or DeepSeek-R1 DeepSeekAI2025DeepSeekR1IR generate and re-rank thousands of reasoning traces per query, achieving state-of-the-art results on, e.g., math or coding quan2025codeelo; hendryckstest2021; rein2024gpqa; eval-harness; open-llm-leaderboard-v2; Clark2018ThinkYH. However, these gains come at a steep cost fan2025cothinktokenefficientreasoninginstruct; kim2025costdynamicreasoningdemystifying. Some studies question whether such models truly learn logical structure or merely exploit surface-level patterns fan2025cothinktokenefficientreasoninginstruct; shojaee2025illusion; xie2024memorization. Curriculum learning has been shown to enhance training robustness and generalization bengio2009curriculum; bursztyn2022compositional, yet no prior framework offers a flexible framework for task synthesis for automatic curriculum generation with symbolic evaluation for reasoning at scale. SLR addresses this gap.
@@ 3.1 Task Specification (Input)
Language Specification ($\mathcal{L}$):
We define a language $\mathcal{L}=(\mathcal{V},\mathcal{G})$ that specifies the building blocks for task generation. The Vocabulary $\mathcal{V}$ comprises a set of constant, function, and predicate symbols that form the syntax for generating rules, examples, and background knowledge. The vocabulary induces the Herbrand base $\mathrm{HB}(\mathcal{V})$, which is the set of all syntactically valid ground atoms (facts) lloyd2012foundations. The Grammar $\mathcal{G}$ is formed by a set of semantic rules that filter the Herbrand base to include only meaningful atoms, $\mathrm{HB}_{\mathcal{G}}(\mathcal{V})$. For instance, a color can be assigned to a car (car_color(car, red)) but not to semantically incompatible objects.
Task Configuration ($\Theta$): The configuration parameters $\Theta=\langle R_{\text{sample}},R_{\text{len}},B_{\pi},\kappa\rangle$ give control over the synthesis process. The (i) Rule Sampling Policy ($R_{\text{sample}})$ controls the synthesis of the ground truth rule $R^{\star}$, which can either be sampled randomly (Uniform Sampling) or generated via an LLM (LLM-Guided Generation). To ensure the LLM produces diverse and challenging logic rules, we leverage an exhaustive prompt (see App. E) that covers a wide array of logical structures and Prolog features (containing arithmetics, recursions, variables, cuts, or comparison operators, etc.). The (ii) Rule Length ($R_{\text{len}}$) specifies the number of literals in the body of the ground-truth rule $R^{\star}$. The (iii) Background Sampling Policy ($B_{\pi}$) defines a probability mass function that assigns a selection probability to each ground atom in the grammar-filtered Herbrand base $\mathrm{HB}_{\mathcal{G}}(\mathcal{V})$, enabling designers to encode priors on the data distribution (e.g., uniform). We also include mirror sampling, where backgrounds for $(E^{+},E^{-})$ are identical except for ground atoms relevant to $R^{\star}$. The (vi) Problem Size ($\kappa=(\kappa_{\text{pos}},\kappa_{\text{neg}})$) specifies the target number of positive ($\kappa_{\text{pos}}=|E^{+}|$) and negative ($\kappa_{\text{neg}}=|E^{-}|$) examples. This directly controls the size and class balance of each generated task.
@@ 3.2 Task Synthesis (Generation)
The task synthesizer (Fig. 1, center) is an automated process detailed in Alg. 1. Given a high-level task specification, the synthesizer generates complete and solvable ILP problems. The process consists of two main phases: rule synthesis and background synthesis.
@@
(ii) Assign Label: The function determines whether a query atom $q$ (that is the ground atom of the target predicate $h$) is logically entailed by the sampled background $b$ and the ground-truth rule $R^{\star}$ (i.e., whether $b\cup R^{\star}\models q$ holds). This produces a label (positive or negative) for the query. We denote the labeling function as:
$\textsc{Assign}(R^{\star},b)=\begin{cases}(1,q),&\text{if}~b\cup R^{\star}\models q\\ (0,q),&\text{otherwise}\end{cases}$
@@ 3.3 Training and Evaluation
Verifiable Logic Rewards.
The SymbolicJudge is a core component of SLR, used for both training and evaluation. It deterministically assesses candidate hypotheses for logic tasks by executing them against a validation program and providing verifiable logic rewards (see Fig. 2). Specifically, it checks whether all positive examples ($E^{+}$) are entailed and all negative examples ($E^{-}$) are not entailed, producing three types of feedback:
Syntax Score: A binary score (0 or 1) that indicates whether the candidate hypothesis $H$ is a syntactically valid Prolog rule. If $H$ is invalid, subsequent scores are set to 0.
OverallScore: Indicates perfect task completion. It returns 1 if the hypothesis $H$, in conjunction with the background knowledge $B$ entails all positive examples ($E^{+}$) and refutes all negative examples ($E^{-}$). Otherwise, the score is 0.
[…]
PartialScore: The proportion of examples (from $E^{+}\cup E^{-}$) that are correctly classified by $H$ given $B$, providing a continuous measure of progress even if the task is not perfectly solved.
The numerator counts correctly entailed positives and refuted negatives, and the denominator normalizes the score to $\llbracket 0,1\rrbracket$.
Model Evaluation.
SLR streamlines the creation of logical reasoning benchmark datasets for systematic model evaluation. By specifying various combinations of task language and configuration, users can generate diverse ILP tasks that span a broad range of domains, prompt styles, and reasoning complexities. Each synthesized task comprises a natural language prompt, the corresponding ground-truth rule, and an executable validation program for assessment by the Symbolic Judge.
Model Training.
SLR enables automated training loops, with model feedback available in two flavors. For supervised fine-tuning (SFT), the ground-truth rule $R^{\star}$ serves as the training target, allowing for cross-entropy loss updates based on predicted rules. For reinforcement learning, the symbolic judge provides verifiable rewards (RLVR) to guide the model’s policy updates. This cohesive, automated pipeline enables scalable generation, evaluation, and training of LLMs, facilitating systematic advances in logical reasoning capabilities.
width=⟧figs/teaser.pdf
@@
Novelty and Systematic Generalization.
As SLR is fully synthetic, overlap with existing data is statistically negligible, making it robust against data leakage and memorization. By default, test set ground-truth rules ($R^{\star}$) are excluded from the training set, guaranteeing that inputs and outputs at test time are completely novel to the model. This enables assessing whether the model is capable of systematically generalizing to entirely new rules.

### v5

@@ Abstract
Lukas Helff1,2, Ahmad Omar1, Felix Friedrich3thanks: work done while at TU Darmstadt/hessian.AI/Lab1141, Antonia Wüst1, Hikaru Shindo1, Tim Woydt1, Rupert Mitchell1,2, Patrick Schramowski1,2,4,5, Wolfgang Stammer1,2,6, Kristian Kersting1,2,4,6,7 1TU Darmstadt 2hessian.AI 3Meta FAIR 4DFKI 5CERTAIN, Germany 6Lab1141 7Centre for Cognitive Science, Darmstadt Code: https://github.com/ml-research/ScalableLogicalReasoning Data: https://huggingface.co/datasets/AIML-TUDA/SLR-Bench
@@ 1 Introduction
To tackle these challenges, we introduce SLR (Scalable Logical Reasoning), an open-source framework for evaluating and training models in inductive logical reasoning. Given a user-defined logic task (Fig. 1, left), the task synthesizer (center) automatically generates novel reasoning tasks of controllable complexity. Each task comes with (i) a latent ground-truth rule, (ii) an executable validation program, and (iii) an instruction prompt. The ground-truth rule serves as the reference answer, while the validation program deterministically evaluates any candidate hypothesis. SLR supports both systematic model evaluation (Fig. 1, top right) and downstream model training, via supervised finetuning or reinforcement learning with rewards provided by the integrated symbolic judge (Fig. 1, bottom right). SLR ’s fully symbolic and automated pipeline removes the need for human annotation and prevents dataset overlap.
Leveraging SLR, we present SLR-Bench (Fig. 3), a “19k task benchmark” that forms a twenty-level curriculum of increasing logical complexity. These levels are further organized into four curriculum tiers: basic, easy, medium, and hard. Each task is unique, with a precise and systematic assessment of inductive logical reasoning skills. In evaluations, we find that while LLMs are generally well-versed in generating valid rules, robust logical reasoning remains challenging. Performance declines sharply as task complexity increases. Scaling model size brings only marginal improvements, while scaling test-time compute boosts reasoning, but returns diminish as complexity rises.
Beyond benchmarking, SLR enables curriculum learning, boosting reasoning both in-domain and across established reasoning benchmarks. SLR-tuned models not only surpasses conventional LLMs on SLR-Bench, but also outperforms reasoning LLMs, such as Gemini-2.0-flash-thinking, while using fewer inference tokens. Notably, these enhanced reasoning capabilities generalize downstream (e.g., GPQA rein2024gpqa, and CLUTRR sinha2019clutrr).
@@
2 Related Work
⟦t!⟧ Reasoning Data Evaluation Task Custom Curriculum Scalable Dataset Type Creation Methodology Synthesis Tasks Learning Complexity LogiQA liu2020logiqa Deduction Human MC FOLIO han2022folio Deduction Human EM AbductionRules young2022abductionrules Abduction Synthetic/Human EM FineLogic zhou2025FineLogic Deduction DS collection Symbolic/LLM Big-Bench phan2025humanitysexam Mix DS collection Mix LogiGLUE luo2024logiglue Mix DS collection MCQA/Auto CLUTRR sinha2019clutrr Induction Synthetic EM KOR-Bench ma2024korbenchbenchmarkinglanguagemodels Mix Synthetic/Human EM/Symbolic (✓) PrOntoQA PrOntoQA Deduction Synthetic EM SynLogic liu2025synlogic Deduction Synthetic EM FLD morishita202FLD Deduction Synthetic Symbolic (✓) K&K xie2025logicrlunleashingllmreasoning; xie2024memorization Deduction Synthetic EM (✓) (✓) ZebraLogic zebralogic2025 Deduction Synthetic EM (✓) (✓) SLR (ours) Induction Synthetic Symbolic
@@ 2 Related Work
Evaluating LLMs’ Logical Reasoning. Tab. 1 provides an overview of existing logical reasoning benchmarks in terms of inference types, dataset origins, and evaluation formats. Notable datasets include LogiQA/2.0 liu2020logiqa; Hanmeng2023logicqa2, FOLIO han2022folio (deductive reasoning), AbductionRules young2022abductionrules (abductive reasoning), bAbI Weston2015TowardsAQ, and CLUTRR sinha2019clutrr (synthetic QA with inductive reasoning). Aggregate testbeds such as BIG-Bench kazemi2025bigbenchextrahard; suzgun2023bigbench, HLE phan2025humanitysexam, FineLogic zhou2025FineLogic, and LogiGLUE luo2024logiglue span a range of tasks and inference styles. Other benchmarks, such as Proofwriter, PrOntoQA, FLD, Multi-LogiEval, SynLogic, ZebraLogic, and the K&K Sandbox tafjord2021proofwriter; PrOntoQA; morishita202FLD; patel2024multilogieval; liu2025synlogic; zebralogic2025; xie2025logicrlunleashingllmreasoning; xie2024memorization generate tasks from fixed ontologies, often with exact-match evaluation. Classic ILP datasets (e.g. Mutagenesis debnath1991mutagenesis) lack scalability and natural-language integration. In contrast, SLR introduces scalable, curriculum-based synthesis with controllable difficulty and verifiable evaluation, addressing key gaps in prior works.
Limits and Promises of Reasoning LLMs. LLMs like GPT-4 openai2023gpt4, Llama-3 grattafiori2024llama3herdmodels, and Qwen bai2023qwen can handle basic reasoning and coding tasks but often struggle with true abstraction shojaee2025illusion; xie2024memorization. Recent reasoning LLMs attempt to bridge this gap by scaling test-time compute. Systems like OpenAI’s o1/o3 openai_o3_system_card_2025 or DeepSeek-R1 DeepSeekAI2025DeepSeekR1IR generate and re-rank thousands of reasoning traces per query, achieving state-of-the-art results on, e.g., math or coding quan2025codeelo; hendryckstest2021; rein2024gpqa; eval-harness. However, these gains come at a steep cost fan2025cothinktokenefficientreasoninginstruct; kim2025costdynamicreasoningdemystifying. Some studies question whether such models truly learn logical structure or merely exploit surface-level patterns fan2025cothinktokenefficientreasoninginstruct; shojaee2025illusion; xie2024memorization. Curriculum learning has been shown to enhance training robustness and generalization bengio2009curriculum; bursztyn2022compositional, yet no prior framework offers a flexible framework for task synthesis for automatic curriculum generation with symbolic evaluation for reasoning at scale. SLR addresses this gap.
@@ 3 SLR: Automatic Benchmark Synthesis
Task Notation. SLR follows the Learning-from-Entailment (LFE) paradigm (Raedt1997LogicalSF) in ILP, where each task is defined as $\mathcal{I}=(B,E^{+},E^{-})$ consisting of background knowledge $B$, positive examples $E^{+}$, and negative examples $E^{-}$. A candidate hypothesis $H$ solves the task if and only if $B\cup H\models E^{+}$ and $B\cup H\not\models E^{-}$. Logical entailment ($\models$) is evaluated via Prolog execution under the closed-world assumption (details see App. A).
Motivating Example. Consider a simple train domain where each train consists of cars described by attributes like color, length, or roof type. The learning goal is to induce a rule, e.g., “the train has a yellow car.” Here, the background knowledge ($B$) contains all known facts about each train, e.g., which cars it includes and their attributes. The positive examples ($E^{+}$) all have yellow cars, while the negative examples ($E^{-}$) do not have any yellow cars. During synthesis, SLR automatically generates such tasks by varying both the background knowledge and the ground-truth rule, effectively scaling the complexity of the reasoning problem.
@@ 3.1 Task Specification (Input)
Language Specification ($\mathcal{L}$): We define a language $\mathcal{L}=(\mathcal{V},\mathcal{G})$ that specifies the building blocks for task generation. The Vocabulary $\mathcal{V}$ comprises a set of constant, function, and predicate symbols that form the syntax for generating rules, examples, and background knowledge. The vocabulary induces the Herbrand base $\mathrm{HB}(\mathcal{V})$, which is the set of all syntactically valid ground atoms (facts) lloyd2012foundations. The Grammar $\mathcal{G}$ is formed by a set of semantic rules that filter the Herbrand base to include only meaningful atoms, $\mathrm{HB}_{\mathcal{G}}(\mathcal{V})$. For instance, a color can be assigned to a car (car_color(car, red)) but not to semantically incompatible objects.
@@
Task Configuration ($\Theta$): The configuration parameters $\Theta=\langle R_{\text{sample}},R_{\text{len}},B_{\pi},\kappa\rangle$ give control over the synthesis process. The (i) Rule Sampling Policy ($R_{\text{sample}})$ controls the synthesis of the ground truth rule $R^{\star}$, which can either be sampled randomly (Uniform Sampling) or generated via an LLM (LLM-Guided Generation). To ensure the LLM produces diverse and challenging logic rules, we leverage an exhaustive prompt (see App. B.5) that covers a wide array of logical structures and Prolog features (containing arithmetics, recursions, variables, cuts, or comparison operators, etc.). The (ii) Rule Length ($R_{\text{len}}$) specifies the number of literals in the body of the ground-truth rule $R^{\star}$. Structural predicates such as has_car(T,C), which merely link entities and impose no constraints, are not counted. The (iii) Background Sampling Policy ($B_{\pi}$) defines a probability mass function that assigns a selection probability to each ground atom in the grammar-filtered Herbrand base $\mathrm{HB}_{\mathcal{G}}(\mathcal{V})$, enabling designers to encode priors on the data distribution (e.g., uniform). We further introduce mirror sampling, where backgrounds for $(E^{+},E^{-})$ are identical except for atoms relevant to $R^{\star}$. For each positive example, a corresponding “mirror” negative is created by randomly altering the rule-specific atoms. The (vi) Problem Size ($\kappa=(\kappa_{\text{pos}},\kappa_{\text{neg}})$) specifies the target number of positive ($|E^{+}|$) and negative ($|E^{-}|$) examples. This directly controls the size and class balance of each generated task.
@@ 3.2 Task Synthesis (Generation)
The task synthesizer (Fig. 1, center) is an automated process detailed in Alg. 1. Given a task specification, the synthesizer generates complete and solvable ILP problems. It comprises two main phases: rule synthesis and background synthesis.
width=.99⟧figs/judge/judge.pdf
@@
(ii) Assign Label: The function determines whether a query atom $q$ (that is the ground atom of the target predicate $h$) is logically entailed by the sampled background $b$ and the ground-truth rule $R^{\star}$ (i.e., whether $b\cup R^{\star}\models q$ holds). This produces a label (positive or negative) for the query. We denote the labeling function as: $\textsc{AssignLabel}(R^{\star},b)=(1,q)$ if $b\cup R^{\star}\models q$, and negative otherwise, $(0,q)$.
@@ 3.3 Training and Evaluation
Verifiable Logic Rewards. In SLR verifiable rewards are implemented via the Symbolic judge (see Fig. 2), which is used for both training and evaluation. It deterministically assesses candidate hypotheses for logic tasks by executing them against a validation program and providing rewards signals. Specifically, it checks whether all positive examples ($E^{+}$) are entailed and all negative examples ($E^{-}$) are not entailed (see App. B.4).
width=.99⟧figs/teaser.pdf
@@
Model Evaluation. SLR streamlines the creation of logical reasoning benchmark datasets for systematic model evaluation. By specifying various combinations of task language and configuration, users can automatically generate diverse tasks that span a broad range of domains, prompt styles, and reasoning complexities. Each synthesized task comprises a natural language prompt, a ground-truth rule, and an executable validation program.
Model Training. SLR enables automated training loops, with two types of model feedback. In supervised fine-tuning (SFT), the ground-truth rule $R^{\star}$ acts as the training target, enabling gradient-based updates from predicted rules. In reinforcement learning, the verifiable logic rewards provide a reward signal (RLVR) that guides policy updates. This cohesive pipeline enables scalable generation, evaluation, and training of LLMs, driving systematic advances in logical reasoning capabilities.
@@
Novelty and Systematic Generalization. The generator enforces a strict separation between training and test splits, ensuring that ground-truth rules ($R^{\star}$) and background knowledge ($B$) are entirely distinct. As SLR is fully synthetic, overlap with existing data is statistically negligible, making it robust against data leakage and memorization. This enables assessing whether the model is capable of systematically generalizing to entirely new rules.